Shift Summary Eve Wednesday 03/Nov/99

Herman B. White (hwhite@fnal.gov)
Thu, 4 Nov 1999 00:08:01 -0600


Shift Summary for Wednesday Evening, 03/Nov/99

On Shift: A. Alavi-Harati, H. White

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Shutdown and Startup

========================================================================
Accesses during shift: Hall was opened until 18:00

========================================================================
Overview of shift:
Exercising components before beam arrives.

13:30 Previous hour Laser scan #14897 complete.
YBH starts test of multiple start/stops runs with COSMICs to test
KTevana errors and Kfrends. Some DYC busy problems, requiring a
one time hardware reset at the KQ crates. See Log XXII p.150
for details. Test successful after that and complete 14:00.
14:20 Etot LV discovered tripped off. Run 14901 executes properly.
15:00 Cosmic run 14902 successfully writes to tape.
16:-- J. Ladue continues to work on kumquat hardware.
17:30 Operations Crew start interlocking NM3 and NM4.
18:00 Interlocks complete, and beam expected in two hours.
18:40 Gas Checklist complete
19:18 NM4AN power supplies unlocked and turned on. Magnet power supply
cycled and polarity changed in preparation for muon run.
20:00 New update: Establishing beam in Tevatron
Estimate 2 hours to FT beam.
23:23 Tuning to the SY dump, and recovering from F1 quench


Notes: A known problem exists in station 19 RR191. The CAMAC crate will
be fixed or replaced possibly tomorrow. The Slow DAQ alarm should
be noted accordingly.

Addition tapes labeled..

Next Shift Information:

Rick Tesarek special request: Please document all KTevana
errors (during beam running) with the following information:

Run#, Spill#, Error type, Stream, RWP(crate)